# Neuroplastic Expansion in Deep Reinforcement Learning

### v1

@@ 4.2 Method: Neuroplastic Expansion
Hill et al. (2010) demonstrated that the cerebral cortex undergoes expansion in response to environmental stimuli, learning, and novel experiences. This dynamic growth and reorganization enhances cognitive and decision-making abilities, fostering adaptability to changing situations. Motivated by our preliminary results and inspired by this biological mechanism, we present Neuroplastic Expansion (NE), which trains RL agents with a dynamic expansion network and seeks to maintain the ability to retain learning capacity and adaptability throughout the entire learning cycle.
Elastic Neuron Generation
@@
While incremental random topology growth has contributed to maintaining plasticity, we aim to elucidate the underlying mechanisms that drive its effectiveness. We pursue this goal by delving into the dormant neuron theory (Sokar et al., 2023), a prevalent explanation for plasticity loss currently. This theory posits that neurons within a network become inactive during training, generally measured by their activations going to zero (Lu et al., 2019). This hinders said neurons’ ability to learn, reducing the network’s overall ability to efficiently process new situations and the outputs gradually diminish, ultimately dropping below the activation function threshold (Lu et al., 2019). Meanwhile, another line of work revealed that parameters lacking prompt exposure to high-gradient stimuli are prone to slipping into dormancy. Drawing from the aforementioned findings, we derive the following observation:
@@
The naive random topology growth sporadically introduces new connections that can generate significant gradients for neurons nearing dormancy, thereby supplying them with potent signals during backpropagation and continue learning.
If the above conjecture aligns with the underlying principles in deep RL, choosing topology candidates based on the gradient magnitude could alleviate plasticity loss more effectively than randomly expanding the network connections. To this end, we follow the sparse network training framework in Tan et al. (2022). Prior to each topology expansion, a batch of transitions is sampled from buffer $D$ to compute the magnitude of gradient $L_{t}$ through the whole model parameters, formalized as $|\nabla_{\theta}L|$. Then $k$ connections capable of yielding the highest magnitude are chosen to augment the topology: $\mathbb{I}_{grow}=Topk_{i\notin\theta^{l}}(|\nabla_{\theta}^{l}L|)$. The topology growing process $G_{k}$ adaptively starts every $\Delta T$ step. And at timestep $t$ the $k$ is decayed in a cosine annealing manner: $\frac{\alpha}{2}(1+\cos(\frac{t\pi}{T_{end}}))$, where $\alpha$ denotes the discount factor and $T_{end}$ is the shut down step.
@@
Figure 3: Performance comparison and plasticity evaluation of random growth, our proposed growth method, and pruning mechanisms (seven independent seeds).
@@
The results depicted in Figure 3 indicate that substituting random growth with the selection of new connections based on potentially provided gradients can enhance the agent’s performance and mitigate the dormant neuron issue. Furthermore, we compare two commonly used topology initial strategies: uniform and Erdos-Rrenyi (Evci et al., 2020), and we find that initializing the network in Erdos-Rrenyi type is slightly more effective for RL. Erdos-Rrenyi enables the number of connections in a sparse layer to scale with the sum of the number of output and input channels, which in turn ensures more stable computations in small networks. Furthermore, the agent can also perform more efficient learning in the early stage.
Algorithm 1 Neuroplastic Expansion TD3
$\pi_{\phi}$: All parameters in actor. $Q_{\theta_{\{1,2\}}}$: All parameters in critics, $M_{l}$: Sparse mask in layer $l$. Stage 1: Neuroplastic Expention (every $\Delta T$)
for each $l_{\phi}\in\pi_{\phi}$, $l_{\theta}\in Q_{\theta_{\{1,2\}}}$ do
     $\mathbb{I}_{grow}^{l},\mathbb{I}_{prune}^{l}\leftarrow G_{k,\omega}.$triangleright Eq.2
     Generate $M_{l_{\phi}},M_{l_{\theta}}\text{ via }\mathbb{I}_{grow},\mathbb{I}_{prune}$
end for
Stage 2: Train the RL policy
     while $t<$ Max training step do
         $a\leftarrow\pi_{\breve{\phi}}(s)$ (with Gaussian noise)
         Observe $r$ and new state $s^{\prime}$
         Fill $D$ with $(s,a,r,s^{\prime})$
         Sample from buffer $D$ via $ER$
         Update $Q_{\breve{\theta}_{\{1,2\}}},\pi_{\breve{\phi}}$ based on TD3
     end while
@@
After replacing the naive growth with our gradient guidance growing process $G_{k}$, a new challenge emerges, i.e., dormant neurons exhibit a stark decline in their representational potency while tenaciously occupying network capacity. Such inactive neurons could be inherent from the initialization or introduced into the network topology through the growth schedule. This situation may lead to a topology that saturates quickly with disproportionately small computational power. Moreover, Lu et al. (2019), Liu et al. (2019) demonstrated that parameters solely forward-linked to dormant neurons are omitted from the gradient calculation process, rendering them ineffective as they fail to receive meaningful guidance signals. We hypothesize that pruning and resetting of dormant neurons following the topology growth could address this issue.
This process aims to free up space for new connections, reactivating dormant neurons as new candidates. According to (Ceron et al., 2024a, b), sparsifying the value-based agents’ network can often enhance the performance, which offers assurance for our real-time pruning empirically. Consequently, we introduce a synchronous dormant neuron prune-and-reset mechanism into the topology growth schedule $G_{k}$.
Here, we adopt the calculation method $f(\cdot)$ showed in Eq.1 and set $\tau$ to be 00, which means only considering fully dormant neurons:$\mathbb{I}{prune}=f(\theta_{i})\leq 0$. To facilitate the topology’s gradual growth, we set the pruning upper bound $\omega$ for each iteration and the prune range for each layer $l$ is $Clip(0,\mathbb{I}_{prune}^{l},\omega\times\mathbb{I}_{grow}^{l})$. Therefore, the final two parameters that control topology change are $k$ and $\omega$ so we will use $G_{k,\omega}$ to refer to our whole grow-prune schedule. The complete growth-pruning schedule for each layer $l$ in actor $\phi$ and critic $\theta$ can be rewritten as
@@
$G_{k,\omega}\left\{\begin{array}[]{l}1.\;\mathbb{I}_{grow}^{l}=Topk_{i\notin\breve{\phi}^{l}}(|\nabla_{\phi}^{l}L_{t}^{\phi}|)+Topk_{i\notin\breve{\theta}^{l}}(|\nabla_{\theta}^{l}L_{t}^{\theta}|)\\ 2.\;\mathbb{I}_{prune}^{l}=f(\breve{\phi}^{l}_{i}\setminus\mathbb{I}_{grow})\leq 0+f(\breve{\theta}^{l}_{i}\setminus\mathbb{I}_{grow})\leq 0;\end{array}\right.$ (2)
@@
We then use $\mathbb{I}_{grow}^{l},\mathbb{I}_{prune}^{l}$ to generate $\{0,1\}$ masks and dot multiply the network parameters to achieve the evolution of the actor topology $\breve{\phi}$ and critic topology $\breve{\theta}$. The results in Figure 3 show that with the addition of the pruning mechanism after each growth round, the plasticity of the RL network is further preserved and the performance is further improved.
@@
As discussed above, the proposed elastic neuron generation and dormant neuron pruning-based dynamic growing approach can greatly mitigate the plasticity loss. This approach smoothly maintains a higher level of active neurons, leading to improved performance. However, a more careful investigation of the results with different seeds reveals that there is a potential risk of performance oscillation during the later learning stages, as RL is sensitive to random seeds (Eimer et al., 2023). This is summarized in the left part of Figure 5.
This behavior suggests that the frequent topology alternations may inadvertently cause the agent’s network to forget previously acquired knowledge. Right part of Figure 5 illustrates the behavior of the agent during the dynamic topology evolution phase. It is worth noting that the policy necessitates the sequential use of three techniques: back straightening, kneeling, and bouncing off the feet. However, a more frequent altering of the topology could result in the forgetting of knowledge acquired initially, which could result in stability issue.
To address this challenge for achieving a better tradeoff, we propose an effective technique to consolidate the agent’s network based on experience review. We consider the fluctuation rate of dormant neurons $\nabla f(\theta)$ to determine whether it is necessary to review earlier experiences for alleviating catastrophic forgetting. Specifically, $\nabla f(\theta)$ measures the slope of changes of the number of dormant neurons (within the critic network) during the recent $C$ steps, i.e., $[150,450]$. A larger positive value of $\nabla f(\theta)$ corresponds to a relatively greater reduction of dormant neurons from our proposed dynamic growing network, while a small value of $\nabla f(\theta)$ close to zero indicates a bottleneck in improved plasticity gains, which also indicates a limited space for further topology expansion and marks the later stage of policy learning (Figure 5). We generally observe that $\nabla f(\theta)$ decreases during the course of training, as elastic neuron generation and dormant neuron pruning will lead to a larger change in the number of dormant neurons in the early stage (and cannot always enlarge the proportion of active neurons due to limited network capacity). We encourage the agent to review earlier experiences when there are more dormant neurons. At each training step, a random number $m\in[0,1]$ is selected, if $m$ exceeds the dynamic threshold $\epsilon$, the sampling is conducted from the initial quarter of the buffer; otherwise, the conventional sampling method is employed.
@@ 5 Experiments
Figure 4: Instability of Neuroplastic Expansion without neuron consolidation via experience review during the later stage.
Figure 5: How $\epsilon$ changes with plasticity measurement. In practice, we set a lower bound for $\epsilon$ to prevent agents from only reviewing historical experience.
In this section, we conduct comprehensive experiments to understand our proposed Neuroplastic Expansion NE method. We investigate the following key questions: (i) How does NE improve the performance of the backbone deep RL method and compare against previous baselines? (ii) Can NE effectively mitigate loss of plasticity? (iii) What are the effects of NE on the policy and value network and the importance of different components? (iv) Can our method be applied to other deep RL methods with more complex image inputs? (v) Can it adapt in continual RL scenarios?
5.1 Performance Comparison
Experimental Setup
We first conduct a series of experiments based on the standard continuous control tasks on OpenAI Gym (Brockman, 2016) simulated by MuJoCo (Todorov et al., 2012). We compare NE against strong baselines including Reset (Nikishin et al., 2022), ReDo (Sokar et al., 2023), Layer Normalization (LN) (Lyle et al., 2024), and Plasticity Injection (PI) (Nikishin et al., 2024). To mitigate implementation bias and ensure a fair comparison, all baselines are implemented based on official implementations (Sokar et al., 2023, Nikishin et al., 2024, 2022), with the same architecture (See Appendix B.1). We build all baselines based on the popular TD3 (Fujimoto et al., 2018) algorithm (where results based on other backbone deep RL methods are discussed in Section 5.3). For each algorithm we run 7 seeds and report the mean and standard deviation. A detailed description of the hyperparameters and setup can be found in Appendix B.2. The growth termination scale for NE is set to match that of the baselines. For all tasks, NE initializes agents from $25\%$ of total capacity and grows asymptotically.
Results
As shown in Figure 6, Reset inevitably suffers from periodic sharp performance drops after each reset for reviving dormant neurons to a large extent, as it can lead to the loss of reusable knowledge that necessitates relearning and impedes the overall progress. LN and ReDo demonstrate more stable performance compared to Reset, but result in a lower proportion of active neurons, while PI maintains the highest level of neuron activation, but only performs in the relatively easier Hopper environment. Our method achieves significant and consistent performance improvements compared to these strong recent baselines in terms of learning efficiency and final performance, with a larger margin in more complex environments. In addition, it effectively improves the loss of plasticity as it maintains a higher level of active neuron ratio throughout training compared to the most competitive LN method. It also better trade-off between performance and neuron utilization, as it performs stably and achieves better performance than PI, which has the highest neuron activation but suboptimal performance in most tasks. It is also worth noting that our proposed method is orthogonal to existing techniques, where additional experiments for building it upon these established baselines can be found in Appendix C allowing for further improvements.
@@ Results
Figure 6: Performance comparisons on OpenAI MuJoCo environments (Todorov et al., 2012) (a) Hopper (b) Walker2D (c) HalfCheetah (d) Ant (e) Humanoid. (f) Results of the activated neurons ratio. Our method outperforms baselines in all four tasks and maintains high activate neurons ratio.
Ablation Study.

### v3

@@ 3.2 Method: Neuroplastic Expansion
Hill et al. (2010) demonstrated that the cerebral cortex expands in response to environmental stimuli, learning, and novel experiences. This dynamic growth and reorganization enhance cognitive and decision-making abilities, fostering adaptability to changing situations. Motivated by our preliminary results and inspired by this biological mechanism, we present Neuroplastic Expansion (NE), a training framework for RL agents that employs a dynamically expanding network to retain learning capacity and adaptability throughout the entire learning process.
Elastic Topology Generation
@@
While incremental random topology growth has contributed to maintaining plasticity, we aim to elucidate the underlying mechanisms that drive its effectiveness. To this end, we explore the dormant neuron theory (Sokar et al., 2023), a prevalent explanation for plasticity loss. This theory posits that neurons within a network become inactive during training, typically when their activation values approach zero (Lu et al., 2019). When this occurs, these neurons lose their ability to learn, reducing the network’s overall capacity to efficiently process new situations. Their outputs gradually diminish, ultimately falling below the activation function threshold (Lu et al., 2019). Meanwhile, prior research has shown that parameters that lack prompt exposure to high-gradient stimuli are more prone to dormancy. Drawing on these findings, we derive the following observation:
@@
Naive random topology growth sporadically introduces new connections that can generate significant gradients for neurons nearing dormancy, thereby supplying them with strong backpropagation signals and allowing them to continue learning.
@@
Figure 3: Performance comparison and plasticity evaluation of random growth, our proposed growth method, and pruning mechanisms. All experiments were conducted with seven independent seeds.
If the above conjecture holds in deep RL, then selecting topology candidates based on gradient magnitude may mitigate plasticity loss more effectively than randomly expanding network connections. To this end, we follow the sparse network training framework in Tan et al. (2022). Before each topology expansion, we sample a batch of transitions from buffer $D$ to compute the gradient magnitude across all model parameters, formalized as $|\nabla_{\theta}L|$. Then, we select $k$ connections that yield the highest gradient magnitudes to augment the topology: $\mathbb{I}_{grow}=ArgTopk_{i\notin\theta^{l}}(|\nabla_{\theta}^{l}L|)$. The topology growing process $G_{k}$ adaptively starts every $\Delta T$ step. We try several growing schedules to achieve warm topology growth to prevent training instability caused by significant network modifications. We chose cosine annealing through practice (see Appendix F.5 for experiments): at timestep $t$ the $k$ is decayed in a cosine annealing manner: $\frac{\alpha}{2}(1+\cos(\frac{t\pi}{T_{end}}))$, where $\alpha$ denotes the discount factor and $T_{end}$ is the shut down step. Finally, we restrict the weights of the newly included neurons (which are randomly initialized) to a specific range for training stabilization (Nauman et al., 2024) via weight clipping (Elsayed et al., 2024). Following standard practice (Liu et al., 2022), we start applying weight clipping on both the input and output weights of a neuron once it is incorporated into the topology.
@@
The results in Figure 3 indicate that substituting random growth with gradient-based selection of new connections enhances agent performance and mitigates dormant neuron issues. Additionally, we compare two commonly used topology initialization strategies: uniform and Erdos-Rrenyi (Evci et al., 2020), and we find that initializing the network using the Erdos-Rrenyi type is slightly more effective for RL. This is because Erdos-Rrenyi enables the number of connections in a sparse layer to scale with the sum of the input and output channels, leading to more stable computations in smaller networks. Consequently, the agent can achieve more efficient learning in the early stages of training.
@@
After replacing the naive growth strategy with our gradient-guided growth process $G_{k}$, a new challenge emerges, i.e., dormant neurons experience a significant decline in representational potency while continuing to occupy network capacity. These inactive neurons may be inherent from initialization or introduced into the network topology through the growth schedule. This situation can lead to a topology that saturates quickly, resulting in disproportionately low computational power. Moreover, Lu et al. (2019); Liu et al. (2019) demonstrated that parameters exclusively forward-linked to dormant neurons are omitted from the gradient calculation process. As a result, they become ineffective, as they fail to receive meaningful guidance signals. We hypothesize that pruning and resetting dormant neurons following topology expansion can address this issue.
This process aims to free up space for new connections, allowing dormant neurons to be reactivated as new candidates. According to Ceron et al. (2024a; b); Sokar et al. (2025), sparsifying the value-based agents’ network can often enhance performance, which empirically supports our real-time pruning approach. Consequently, we introduce a synchronous prune-and-reset mechanism into the topology growth schedule $G_{k}$. Here, we adopt the calculation method $f(\cdot)$ shown in Eq. 1 and set $\tau$ to be 00, which means only considering fully dormant neurons: $\mathbb{I}_{prune}=\{\text{index}(\theta_{i})|f(\theta_{i})=0\}$. To facilitate the topology’s gradual growth and avoid over-pruning, we employ the Truncate process (that operates on the pruning set $\mathbb{I}^{l}_{prune}$) which drops excess elements randomly from the set when its size exceeds a predefined upper bound. The new truncated set is then used for pruning (Refer to Appendix D.2 for the detailed pseudo-code of our Truncate Process). In practice, we define the pruning upper bound in each layer $l$ as $\omega\times|\mathbb{I}^{l}_{grow}|$, where $|\mathbb{I}^{l}_{grow}|$ represents the number of elements in the current grow set. The parameter $\omega\in[0,1)$ acts as a “discount” factor, moderating the pruning relative to the total network growth in layer $l$. Thus, the two final parameters controlling topology evolution are $k$ and $\omega$. We denote our grow-prune schedule as $G_{k,\omega}$. The complete growth-pruning schedule for each layer $l$ in actor $\phi$ and critic $\theta$ networks is defined as:
@@
$G_{k,\omega}=\left\{\begin{array}[]{l}1.\;\mathbb{I}_{grow}^{l}=\text{ArgTopk}_{i\notin\breve{\phi}^{l}}(|\nabla_{\phi}^{l}L_{t}^{\phi}|)\cup\text{ArgTopk}_{i\notin\breve{\theta}^{l}}(|\nabla_{\theta}^{l}L_{t}^{\theta}|)\\ 2.\;\mathbb{I}_{prune}^{l}=\big{\{}\text{Index}(\breve{\phi}^{l}_{i})|f(\breve{\phi}^{l}_{i})=0\big{\}}\cup\big{\{}\text{Index}(\breve{\theta}^{l}_{i})|f(\breve{\theta}^{l}_{i})=0\big{\}}\end{array}\right.$ (2)
@@
We then use $\mathbb{I}_{grow}^{l},\mathbb{I}_{prune}^{l}$ to generate $\{0,1\}$, which are dot-multiplied with the network parameters to achieve the evolution of the actor topology $\breve{\phi}$ and critic topology $\breve{\theta}$. The results in Figure 3 demonstrate that introducing the pruning mechanism after each growth round further preserves the plasticity of the RL network and enhances overall performance. NE is inspired by plasticity injection (PI) (Nikishin et al., 2024). While both NE and PI maintain plasticity by extending networks, PI focuses on one-time large-scale network growth, whereas our method achieves continuous real-time adaptation.
Notably, our sparse network training framework for maintaining stable learning is summarized as follows: (i) Inspired by Evci et al. (2020), we use Erdos-Rrenyi initialization to ensure that the number of connections in a sparse layer scales with the sum of the output and input channels, thereby enhancing initial stability. (ii) We employ cosine annealing to guide gradual network growth, reducing noise caused by network changes. (iii) Following sparse training practices in RL (Tan et al., 2022), we keep the first and last layers dense to ensure the stability of the encoding and decoding.
Figure 4: Instability of NE in the later stage without neuron consolidation via experience review.
@@
As discussed above, the proposed elastic topology generation and dormant neuron pruning-based dynamic growth approach effectively mitigate plasticity loss. This method maintains a higher proportion of active neurons, leading to improved performance. However, a closer examination of the results, i.e. Tab. 8, reveals that NE demonstrates a high standard deviation across seeds in the later stage of training, suggesting a potential risk of performance oscillation near convergence. We hypothesize that frequent network topology changes may introduce slight probabilistic errors in the network structure. Consequently, as the activated neuron ratio decreases, some agents may risk forgetting previously learned skills from the early training stages (e.g., standing in the HalfCheetah task). This phenomenon leads to suboptimal behavior post-convergence and increases variance between seeds.
We examined the behaviors on the HalfCheetah task and found that there are multiple runs in which the agent no longer could stand after a fall (Figure 4). This observation strengthens our confidence in the above conjecture. To address this, we seek a topology-agnostic and simple mechanism to further enhance the stability of NE by mitigating catastrophic forgetting. Meanwhile, prior studies (Rolnick et al., 2019; Zhou et al., 2020) show that experience replay (ER) helps mitigate forgetting and supports network plasticity in both RL and supervised learning. Inspired by these findings, we aim to develop an ER mechanism sensitive to both the training stage and the activated neuron ratio.
In practice, we define the absolute slope of the activated neuron ratio curve as: $\Delta(f(\theta)):=\sum_{i\in\theta}|\Delta f(\theta_{i})|$ as the threshold $\epsilon$. Specifically, $\epsilon$ measures the slope of changes in the number of activated neurons (within the critic network) during the recent $c$ steps, i.e., $[150,450]$. A larger positive value of $\epsilon$ corresponds to a greater reduction in dormant neurons, indicating active plasticity improvements in our dynamically growing network. In contrast, a small $\epsilon$ (close to zero) suggests a bottleneck in plasticity gains, meaning limited capacity for further topology expansion and marking the later stage of policy learning (see Figure 5).
During training, we observe that $\epsilon$ generally decreases over time, since elastic neuron generation and dormant neuron pruning significantly impact the number of dormant neurons in the early training stage. However, due to limited network capacity, they cannot indefinitely increase the proportion of active neurons. To counteract this effect, we encourage the agent to review earlier experiences from the buffer when dormant neurons are more prevalent. At each training step, a random number $m\in[0,1]$ is selected, if $m>\epsilon$, the sampling is conducted from the initial quarter of the buffer; otherwise, the conventional sampling method is employed. We provide pseudocode outlining the full algorithmic process in Appendix D.1.
@@ 4 Experiments
Figure 5: How $\epsilon$ changes with plasticity measurement. In practice, we set a lower bound for $\epsilon$.
In this section, we conduct comprehensive experiments to evaluate whether NE can alleviate the loss of plasticity. We investigate the following key questions: (i4.1)) In a long-term training setting, can NE effectively combat early data overfitting, enhance performance, and attenuate the decline in activation neuron ratio? (ii4.2)) In continuous adaptation task, can NE enable agents to efficiently adapt to new tasks while slowing the dormancy rate of activated neurons, without being constrained by prior learning limitations?
(iii4.3)) What are the effects of NE on the policy and value networks, and what is the importance of its different components? (iv4.4)) Can our method be applied to other deep RL methods with more complex image inputs?
4.1 Long-term Training Tasks
@@
Figure 6: Performance comparisons on OpenAI MuJoCo environments (Todorov et al., 2012): (a) Hopper, (b) Walker2D, (c) HalfCheetah, (d) Ant, (e) Humanoid. (f) Results of the activated neuron ratio. NE outperforms the baselines in four tasks and maintains a high activated neuron ratio.
Experimental Setup
We conduct a series of experiments based on the standard continuous control tasks from OpenAI Gym (Brockman, 2016) simulated by MuJoCo (Todorov et al., 2012) with long-term training setting, i.e. 3M steps $\to$ 6M. We compare NE against strong baselines including Reset (Nikishin et al., 2022), ReDo (Sokar et al., 2023), Layer Normalization (LN) (Lyle et al., 2024b), and Plasticity Injection (PI) (Nikishin et al., 2024). To ensure a fair comparison and mitigate implementation bias, we base all baselines on official implementations (Sokar et al., 2023; Nikishin et al., 2024; 2022), with the same architecture (See Appendix C.1). All baselines are implemented using the TD3 (Fujimoto et al., 2018) algorithm (where results based on other backbone deep RL methods are discussed in Section 4.4). To ensure a fair comparison and reproducibility of the expected performance of all baselines, we follow the recommended hyperparameter setting from Elsayed et al. (2024) (see Appendix C) for all baselines. The clipping parameter $\kappa$ is set to 3. Each algorithm is trained with 7 random seeds, and we report the mean and standard deviation. A detailed description of hyperparameters and experimental setup is in Appendix C.2. For all tasks, NE initializes agents at 25% of total capacity and grows asymptotically. The growth termination scale is aligned with baselines. The score denotes undiscounted episodic return.
Results
As shown in Figure 6, the Reset method suffers from periodic performance drops due to resetting a large portion of layer parameters. This leads to the loss of reusable knowledge, requiring relearning, which impedes overall progress. LN and ReDo demonstrate more stable performance than Reset, but result in a lower proportion of active neurons. PI maintains the highest level of neuron activation, but performs well only in the relatively simpler Hopper environment. NE achieves significant and consistent improvements in learning efficiency and final performance, with a larger margin in more complex environments. Furthermore, NE effectively mitigates plasticity loss, maintaining higher active neuron ratios throughout training compared to the most competitive LN method. Additionally, NE achieves a superior trade-off between performance and neuron utilization. It performs more stably and outperforms PI, which, despite achieving the highest neuron activation, exhibits suboptimal performance in most tasks.
4.2 Continual adaptation
In this section, we investigate the continual ability to learn and adapt to changing environments, which is an essential capability of deep RL agents (Willi* et al., 2024; Elsayed & Mahmood, 2024).
Experimental Setup
[…]
Figure 7: Activated neuron ratio, recorded for three agents with different settings during cyclic training.
Results
The performance of each method in a single environment, when repeatedly learning across a sequence of environments as described above, is summarized in Figure 8. We aim to evaluate (i) the efficiency of the policy in quickly learning a new task after mastering the previous one within a single cycle; (ii) the agent’s ability to retain the benefits of initial learning for the same task across multiple cycles. As shown, Resetting, which involves initializing the parameters after each task, is currently considered the most effective approach.
According to results in Figure 8, NE effectively handles each task without being negatively influenced by prior training. It significantly outperforms the original TD3, reinforcing that standard RL algorithms suffer from catastrophic forgetting in a continual learning setting. We attribute this success to NE’s ability to introduce real-time topology adjustments, which facilitate storing new knowledge and ensuring continuous learning capabilities.
Notably, in certain environments, such as Ant, NE-TD3 retains prior knowledge and seamlessly continues learning. This could be attributed to the experience review module and the preservation of old neurons within the topology, which store critical information essential for successful continual learning. When analyzing column-wise comparisons, NE performs comparably to Reset in most tasks, yet it significantly outperforms Reset in the challenging Humanoid task. The comparison of active neurons for each baseline is shown in Figure 7, which illustrates that even during extended training, NE effectively mitigates neuron dormancy while maintaining efficient learning. In fact, it even outperforms Reset in the early stages of training, despite Reset reinitializing all parameters.
The spike phenomenon in Figure 7 is related to research examining the relationship between input distribution and plasticity. Lu et al. (2019) observed that neurons failing to activate due to the activation function can be reactivated when the input distribution to the network changes. A similar phenomenon has been reported in Ma et al. (2023), where data augmentation was found to alleviate the decline in activated neurons. In the continuous adaptation task we tested, a sudden change in the environment may have altered the input data distribution, leading to a temporary increase in the activated neuron rate. However, after a short training period, the neurons returned to a dormant state.
(a) 1st task: HalfCheetah
(b) 2nd task: Humanoid
(c) 3rd task:Ant
(d) 4th task:Hopper
4.3 Ablation Study